OpenStudy (anonymous):

Can anyone HELP Me out with these two problems PLEASE!!! (x+1)(x^2+5x-2) and the other one is 15x^3y^2+25x^2y^2+5xy^2

OpenStudy (mertsj):

(x+1)(x^2+5x-2)=x^3+5x^2-2x+x^2+5x-2= x^3+6x^2+3x-2

OpenStudy (mertsj):

\[15x ^{3}y ^{2}+25x ^{2}y ^{2}+5xy ^{2}=5xy ^{2}(3x ^{2}+5x+1)\]